# The Schur-Agler class in infinitely many variables

### v1

@@ 1. Introduction
Our goal is to prove that analogues of the Agler decomposition (1.1), the Agler-Pick interpolation theorem (Theorem 1.3), and the transfer function realization (1.2) hold in infinitely many variables. We also wish to make connections to Dirichlet series as in Remark 1.2. Remark 1.1 suggests that we cannot define $f(T)$ in the infinite variable setting using absolutely convergent series. Instead, we shall begin with the simplest definition of the Agler class in infinite variables.
@@
11 11 $=\sum_{j=1}^{\infty}K_{j}(0,0)$
11 11 $=\sum_{j=1}^{\infty}K_{j}(\alpha,0)$
$1-a^{2}$ $=\sum_{j=1}^{\infty}(1-a_{j}^{2})K_{j}(\alpha,\alpha).$
@@
so that
$1\leq\sum_{j=1}^{\infty}K_{j}(\alpha,\alpha).$
On the other hand, we have
@@ 2. Notations and background
is positive semi-definite.
@@
initially defined for vectors indexed by $z\in X$ extends linearly and in a well-defined way to a contractive operator $V$ from $\mathbb{C}\oplus\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*}$ to $\mathbb{C}\oplus\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*}$. We write $V$ in block form
$\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}$
with
  • $A\in\mathbb{C}\cong\mathcal{B}(\mathbb{C},\mathbb{C})$,
  • $B\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*},\mathbb{C})$,
  • $C\in\mathcal{B}(\mathbb{C},\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$,
  • $D\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$,
using the notation $\mathcal{B}(X,Y)$ to denote the bounded linear operators from $X$ to $Y$ (as well as $\mathcal{B}(X)=\mathcal{B}(X,X)$). Let $\Delta(z)\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$ be the diagonal operator sending
@@
$=B(1-\Delta(z)D)^{-1}\Delta(z-w)(1-D\Delta(w))^{-1}C$
@@
Regarding (1) implies (3), by definition, we can make sense of $\tilde{f}(T)$ for $\tilde{f}\in\mathcal{A}_{\infty}$ when $(\|T_{j}\|)_{j\in\mathbb{N}}\in Ball(c_{00})$ and Theorem 1.5 lets us make sense of it when $(\|T_{j}\|)_{j\in\mathbb{N}}\in Ball(\ell^{\infty})$. To prove $\|\tilde{f}(T)\|\leq 1$ when $T$ consists of matrices that are commuting, contractive, simultaneously diagonalizable with joint spectrum $\sigma(T)\subset S$ and eigenspaces of dimension at most $1$, we can give a continuity argument. First, $\|\tilde{f}(rT)\|\leq 1$ holds for $r<1$ because we will have
@@
Let $\delta:X\times X\to\mathbb{C}$ denote the function $\delta(z,w)=1$ if $z=w$ and $\delta(z,w)=0$ if $z\neq w$. There necessarily exist constants $c_{1},c_{2}$ such that for all $n$
$c_{1}\delta\succeq C_{n}\succeq c_{2}\delta.$
This looks more familiar when we view our functions on $X\times X$ as matrices. Since $X\subset\mathbb{D}_{2}^{\infty}$ we can define
$S(z,w)=\prod_{j=1}^{\infty}\frac{1}{1-z_{j}\bar{w}_{j}}\text{ and }S_{j}(z,w)=(1-z_{j}\bar{w}_{j})P(z,w)=\prod_{i\neq j}\frac{1}{1-z_{i}\bar{w}_{i}};$
@@
Using a diagonal argument we can select a subsequence such that each $A_{n,i}$ converges to a positive semi-definite matrix $A_{i}$ as $n\to\infty$. Also, for each $N$ we have
$C_{n}S\succeq\sum_{j=1}^{N}S_{j}A_{n,j}\succeq\sum_{j=1}^{N}A_{n,j}.$
Sending $n\to\infty$ we have
$CS\succeq\sum_{j=1}^{N}S_{j}A_{j}\succeq\sum_{j=1}^{N}A_{j}$
and sending $N\to\infty$ we have
$CS\succeq\sum_{j=1}^{\infty}S_{j}A_{j}\succeq\sum_{j=1}^{\infty}A_{j}$
where the sums converge absolutely. Next, recalling $Z_{j}\otimes\bar{Z}_{j}$ denotes the function $(z,w)\mapsto z_{j}\bar{w}_{j}$ we have

### v2

@@ 1. Introduction
Our goal is to prove that analogues of the Agler decomposition (1.1), the Agler-Pick interpolation theorem (Theorem 1.3), and the transfer function realization (1.2) hold in infinitely many variables. We also wish to make connections to Dirichlet series as in Remark 1.2. Remark 1.1 suggests that we cannot define $f(T)$ in the infinite variable setting using absolutely convergent series.
We would like to remark that this paper is not the first mention of the Agler class in infinitely many variables—see [AL], Section 7. However one of our larger goals is to start with the simplest definition possible and deduce basic properties of the Agler class (such as its functional calculus) as well as to point out some subtleties of the theory. Here is what we imagine to be the simplest definition of the Agler class in infinitely many variables.
@@
$1=\sum_{j=1}^{\infty}K_{j}(0,0),\qquad 1=\sum_{j=1}^{\infty}K_{j}(\alpha,0),\qquad 1-a^{2}=\sum_{j=1}^{\infty}(1-a_{j}^{2})K_{j}(\alpha,\alpha).$
@@
so that $1\leq\sum_{j=1}^{\infty}K_{j}(\alpha,\alpha).$ On the other hand, we have
@@ 2. Notations and background
is positive semi-definite. Here we emphasize that $f$ on $\mathbb{D}_{\rho}$ is defined via $f$’s holomorphic extension from $Ball(c_{00})$ to $Ball(c_{0})$ as in the traditional Schur class of infinitely many variables.
@@
initially defined for vectors indexed by $z\in X$ extends linearly and in a well-defined way to a contractive operator $V$ from $\mathbb{C}\oplus\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*}$ to $\mathbb{C}\oplus\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*}$. We write $V$ in block form $\begin{pmatrix}A&B\\ C&D\\ \end{pmatrix}$ with $A\in\mathbb{C}\cong\mathcal{B}(\mathbb{C},\mathbb{C})$, $B\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*},\mathbb{C})$, $C\in\mathcal{B}(\mathbb{C},\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$, $D\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$, using the notation $\mathcal{B}(X,Y)$ to denote the bounded linear operators from $X$ to $Y$ (as well as $\mathcal{B}(X)=\mathcal{B}(X,X)$). Let $\Delta(z)\in\mathcal{B}(\bigoplus_{j=1}^{\infty}\mathcal{H}_{j}^{*})$ be the diagonal operator sending
@@
$=B(1-\Delta(z)D)^{-1}\Delta(z-w)(1-D\Delta(w))^{-1}C.$
@@
Regarding (1) implies (3), by definition, we can make sense of $\tilde{f}(T)$ for $\tilde{f}\in\mathcal{A}_{\infty}$ when $(\|T_{j}\|)_{j\in\mathbb{N}}\in Ball(c_{00})$ and Theorem 1.5 lets us make sense of it when $(\|T_{j}\|)_{j\in\mathbb{N}}\in Ball(\ell^{\infty})$. To prove $\|\tilde{f}(T)\|\leq 1$ when $T$ consists of matrices that are commuting, contractive, simultaneously diagonalizable with joint spectrum $\sigma(T)\subset X$ and eigenspaces of dimension at most $1$, we can give a continuity argument. First, $\|\tilde{f}(rT)\|\leq 1$ holds for $r<1$ because we will have
@@
Let $\delta:X\times X\to\mathbb{C}$ denote the function $\delta(z,w)=1$ if $z=w$ and $\delta(z,w)=0$ if $z\neq w$. There necessarily exist constants $c_{1},c_{2}$ such that for all $n$, $c_{1}\delta\succeq C_{n}\succeq c_{2}\delta.$ This looks more familiar when we view our functions on $X\times X$ as matrices. Since $X\subset\mathbb{D}_{2}^{\infty}$ we can define
$S(z,w)=\prod_{j=1}^{\infty}\frac{1}{1-z_{j}\bar{w}_{j}}\text{ and }S_{j}(z,w)=(1-z_{j}\bar{w}_{j})S(z,w)=\prod_{i\neq j}\frac{1}{1-z_{i}\bar{w}_{i}};$
@@
Using a diagonal argument we can select a subsequence such that each $A_{n,i}$ converges to a positive semi-definite matrix $A_{i}$ as $n\to\infty$. Also, for each $N$ we have $C_{n}S\succeq\sum_{j=1}^{N}S_{j}A_{n,j}\succeq\sum_{j=1}^{N}A_{n,j}.$ Sending $n\to\infty$ we have $CS\succeq\sum_{j=1}^{N}S_{j}A_{j}\succeq\sum_{j=1}^{N}A_{j}$ and sending $N\to\infty$ we have $CS\succeq\sum_{j=1}^{\infty}S_{j}A_{j}\succeq\sum_{j=1}^{\infty}A_{j}$ where the sums converge absolutely. Next, recalling $Z_{j}\otimes\bar{Z}_{j}$ denotes the function $(z,w)\mapsto z_{j}\bar{w}_{j}$ we have
@@
Acknowledgments
Thank you to the referee for a thoughtful report and for pointing out several typos. Thank you to Professor Daniel Alpay for bringing the reference [AL] to our attention.